Bindley Corporation has a one-year contract to supply motors for all washing machines produced by Rinso Ltd. Rinso manufactures the washers at four locations around the country: New York, Fort Worth, San Diego, and Minneapolis. Plans call for the following numbers of washing machines to be produced at each location:

New York 70,000
Fort Worth 67,000
San Diego 98,000
Minneapolis 85,000


Bindley has three plants that can produce the motors. The plants and production capacities are

Boulder 110,000
Macon 114,000
Gary 121,000

Due to varying production and transportation costs, the profit Bindley earns on each 1,000 units depends on where they were produced and where they were shipped. The following table gives the accounting department estimates of the dollar profit per unit. (Shipment will be made in lots of 1,000.)

SHIPPED TO
PRODUCED AT NEW YORK FORT WORTH SAN DIEGO MINNEAPOLIS
Boulder 15 11 12 13
Macon 22 21 18 14
Gary 25 10 24 19

Question: Find the optimal solution using Microsoft Excel.

Candidate
Solution
Total Shipped
Boulder   
Macon
Gary
Total shipped
Profit
Boulder   
Macon
Gary
Total profit
Answer #1

Let X_{ij} number of lots (each lot is of size 1000) of motors shipped from plant i to location j, where i=1,2,3 correspond to {Boulder, Macon,Gary} and j=1,2,3,4 correspond to {NEW YORK, FORT WORTH, SAN DIEGO,MINNEAPOLIS} respectively.

These are the decision variables.

The total profit (in $1000s) is

15×X11+11 × X12 + 12 × X13 + 13 × X14+ 22 × X21 +21 × X22 + 18 × X23 + 14 × X24+ 25 × X31 + 10 × X32 + 24 × X33 + 19 × X34

The constraints are

Demand at 4 locations

  • Demand at New York is 70,000 units (70 lots). The lots shipped to New York should be equal to 70
    • X11 + X21 + X3170
  • Demand at Fort Worth is 67,000 units (67 lots). The lots shipped to Fort Worth should be equal to 67 \begin{align*} X_{12}+ X_{22}+ X_{32}=67 \end{align*}
  • Demand at San Diego is 98,000 units (98 lots). The lots shipped to San Diego should be equal to 98
    • X13 X23 X33 98
  • Demand at Minneapolis is 85,000 units (85 lots). The lots shipped to Minneapolis should be equal to 85
    • \begin{align*} X_{14}+ X_{24}+ X_{34}=85 \end{align*}

Capacity constraint

  • Capacity of plant at Boulder is 110,000. The total lots shipped from this plant cannot exceed 110
    • \begin{align*} X_{11}+ X_{12}+ X_{13}+ X_{14}\le 110 \end{align*}
  • Capacity of plant at Macon is 114,000. The total lots shipped from this plant cannot exceed 114
    • \begin{align*} X_{21}+ X_{22}+ X_{23}+ X_{24}\le 114 \end{align*}
  • Capacity of plant at Gary is 121,000. The total lots shipped from this plant cannot exceed 121
    • X31 + X32+ X33 +X34 1 21

Shipments are made in lots of 1000s. Since we cannot ship a fraction of lot, Xs are integers.

The LP model is

Maximize

15×X11+11 × X12 + 12 × X13 + 13 × X14+ 22 × X21 +21 × X22 + 18 × X23 + 14 × X24+ 25 × X31 + 10 × X32 + 24 × X33 + 19 × X34

s.t.

  • X11 + X21 + X3170
  • \begin{align*} X_{12}+ X_{22}+ X_{32}=67 \end{align*}
  • X13 X23 X33 98
  • \begin{align*} X_{14}+ X_{24}+ X_{34}=85 \end{align*}
  • \begin{align*} X_{11}+ X_{12}+ X_{13}+ X_{14}\le 110 \end{align*}
  • \begin{align*} X_{21}+ X_{22}+ X_{23}+ X_{24}\le 114 \end{align*}
  • X31 + X32+ X33 +X34 1 21
  • Xij 2 0, Xij - integers, i-1,2,3;j-1,2, 3, 4

Prepare the following sheet

1 Profit per unit SHIPPED TO 3 PRODUCED AT NEW YORK FORT WORTH SAN DIEGO MINNEAPOLIS Capacity 4 Boulder 15 13 12 110000 5 Mac

get this

Profit per unit 1 SHIPPED TO 3 PRODUCED AT | NEW YORK | FORT WORTH SAN DIEGO MINNEAPOLIS Capacity 11 S 13 4 Boulder 12 S 110,

setup the solver using data--->solver

Solver Parameters 1 Profit per unit SHIPPED TO Set Objective: SBS 1 3 PRODUCED AT | NEW YORK | FORT WORTH SAN DIEGO MINNEAPOL

get this

G HH 1 Profit per unit SHIPPED TO 3 PRODUCED AT | NEW YORK | FORT WORTH | SAN DIEGO | MINNEAPOLIS | Capacity 15 $ 12 $ 4 Boul

ans:

SHIPPED TO
PRODUCED AT NEW YORK FORT WORTH SAN DIEGO MINNEAPOLIS Total Shipped (1000s)
Boulder 0 0 0 85 85
Macon 47 67 0 0 114
Gary 23 0 98 0 121
Total shipped 320
Profit NEW YORK FORT WORTH SAN DIEGO MINNEAPOLIS Total Profit ($1000s)
Boulder $           -   $               -   $           -   $          1,105 $       1,105
Macon $      1,034 $          1,407 $           -   $               -   $       2,441
Gary $         575 $               -   $      2,352 $               -   $       2,927
Total profit $       6,473